A body is mass m is moving in a straight line with an initial velocity of 2 m/s from its origin at t=0 and acceleration 6t
1
Expert's answer
2021-08-31T01:49:57-0400

We can find the time moved by the moving body,

"Mass = m\\
V_i= 2 m/s\\
V_f=0 m/s\\
a= 6t\\
F= ma\\
m= \frac{F}{a}\\
m= \frac{F}{\frac{2}{t}}\\
But \space a=\frac{2}{t}\\
\implies 6t = \frac{2}{t}\\
\therefore t=0.57735"